### v5

@@ 1 Introduction
Recent years have witnessed a heated debate in the macro and international finance literature on the extent to which global financial conditions influence national ones and, within that, how far capital controls and flexible exchange rates can insulate national economies from global financial shocks. Following the seminal work of Rey (2015), as well as Borio (2014), Bruno and Shin (2013) and others, there has been widespread consensus that global financial conditions—epitomized by the concept of a global financial cycle—are a key determinant of domestic financial cycles and risk assessments by economic agents.
However, the jury is still out on the extent to which domestic policy instruments can insulate the domestic economy from the global financial cycle. Miranda-Agrippino and Rey (2020) and Passari and Rey (2015), using higher frequency data spanning a wider range of financial instruments over a long period that includes more than 30 years, have brought into question the effectiveness of domestic monetary policies and exchange rate flexibility as macroeconomic insulation devices, as advocated by Mundell (1963) in his famous depiction of the trilemma between exchange rate stability, monetary policy independence and free capital mobility. Yet, several papers, including Aizenman et al. (2008), Aizenman et al. (2016), Obstfeld et al. (2019), Obstfeld (2021) and Shambaugh (2004), find evidence that exchange rate flexibility can be effective in insulating domestic interest rates from the global financial cycle, at least partially. In addition, and over and above exchange rate flexibility, various recent studies find that financial account restrictions and some macro-prudential regulations can also be effective insulation devices (e.g., Akinci and Olmstead-Rumsey, 2018; Buch and Goldberg, 2018; Cerutti et al., 2018; Fendoğlu, 2017; Qureshi et al., 2011). These latter findings indicate that the Mundellian trillema remains, by and large, a suitable characterization of main trade-offs faced by domestic macroeconomic policy in a globalized world.
@@
As we discuss below, two important advantages of extracting principal components from residuals obtained by CCE are to allow (i) for the presence of unobserved or hard-to-measure global factors as well as (ii) for some cross-country heterogeneity in bank responses, which one might reasonably expect and which impart biases in the standard fixed effects estimator. We then establish how the choice of this approach influences our results relative to more standard panel fixed effects regressions in earlier studies. The end-goal is to establish the extent to which the estimated insulation role granted by exchange rate flexibility is sensitive to econometric estimation (or model) uncertainty.
@@
Our main findings are as follows. First, fluctuations in the ratio of non-core to core funding are highly persistent and overwhelmingly determined by global factors. Second, the combination of high cross-country dependence of non-core to core ratios and considerable heterogeneity in which global factors affect the banking system in each country, implies that the standard two-way fixed effects estimator (including country and time fixed effects) does not eliminate the cross-sectional dependence of the residuals, therefore producing likely inconsistent estimates of the response of bank ratios to the various country-specific and global disturbances. This indicates that the estimation methodology that we propose—combining the Pesaran (2006) CCE estimator to extract the relevant global factors (principal components) and then using the mean group estimator to tease out any differences between countries on a floating vs. flexible exchange rate regime—should produce superior results. This estimation methodology then delivers our central result, namely that countries on a fixed exchange rate regime are more susceptible to have global factors affecting their bank funding ratios, or in other words exchange rate flexibility helps insulate the non-core to core ratio from such global factors. We also find that countries with tighter macroprudential policy stances are less affected by the global financial cycle, and this insulation is most effective for countries with a fixed exchange rate. We finally take a closer look at the mechanisms through which the principal components affect banks’ non-core ratios. To this end, we first relate the principal components with the set of observed common factors, and find that most of the variation in banks’ non-core ratios can be explained by changes in the world macroprudential policy stance, the US short-term interest rate and financial crisis effects. In contrast, the VIX, oil prices, and the US real exchange rate generally have a smaller impact on banks’ non-core ratios. Then, using BIS Locational Banking Statistic data, we break down total foreign borrowing by banks into (i) domestic- vs. foreign-currency borrowing, as well as (ii) liabilities provided by foreign banks vs. foreign non-bank financials vs. other foreign sectors. In these specifications, we find that our results are driven mostly by banks’ interbank borrowing—a funding source that is known to be more flighty. The currency denomination does not seem to matter a lot.
@@ 2 Data
Our empirical analysis is based on a data set that combines two critical banking sector balance sheet variables from the International Financial Statistics at quarterly frequency over the period 2004:Q1-2022:Q1 with different domestic macroeconomic controls and global variables. Our sample includes the 31 advanced economies listed in Table 8, [^5] In most specifications, we drop Iceland from our sample as it is an outlier for most variables. Its inclusion reduces the significance of our estimates, as we document in the robustness checks of Section 4.2. producing a balanced panel spanning more than 2,000 country-quarter observations. In this section, we give an overview about the data, with Appendix A providing all of the corresponding data details and sources.
@@ 2.1 Outcome Variables
Consistent with this evidence, we measure non-core funding ratios as the share of foreign liabilities over total deposits (IFS line 26c divided by IFS lines 24 and 25). To show how such a non-core ratio relates to other key balance sheet indicators highlighted in previous work, such as the ratio of bank loans-to-deposits (see Edwards and Vegh (1997), Catão and Rodriguez (2000), Obstfeld et al. (2005, 2019)), we define loan-to-deposit ratio as IFS line 22d divided by IFS lines 24 and 25. As Table 1 shows, the average non-core ratio in our sample is equal to 78% and loan-to-deposit ratios have a mean of 122%. The variations are substantial with standard deviations of about 94% and 58%, respectively.
@@
Figure 1 shows that advanced economies’ non-core funding ratios are positively correlated ($\rho=0.88$) with swings in in banks’ loan-to-deposit ratios, revealing the important role of foreign borrowing in expanding lending beyond banks’ core funding sources (i.e., deposits). Figure 2 in turn shows that the non-core ratio is in general higher for countries on a fixed exchange rate. [^6] Notice that the floaters are in the mean on the left scale that is steadily two thirds lower than its span. Besides this level difference, it also shows conspicuous long-term trend shifts and considerable co-movement between the two country groups. In particular, non-core ratios increased significantly before the Global Financial Crisis, reaching a peak in 2008 where the global non-core ratio rose to 70%, to then decrease to a level of 30% in 2022. Only between 2015-20, there was some short-term divergence in the trends of fixers and floaters.
In the second part of the paper, we examine in more detail the mechanisms through which banks’ non-core funding ratios change. To this end, we employ data on banks’ cross-border liabilities from the BIS Locational Banking Statistics, which we also scale by total deposits. The data have poorer country coverage than our IFS data, but have the advantage of allowing for a breakdown into domestic- vs. foreign-currency liabilities, as well as into foreign liabilities provided by banks, nonbank financials, and all other sectors. The correlation between our benchmark non-core ratio and the overall BIS cross-border liability variable is 88%. Looking further into the different breakdowns of the BIS variable, we find that the correlation coefficient between the non-core ratio and the foreign-currency variable stands at 79%, whereas it is 81% for the domestic-currency variable, indicating that the currency denomination of foreign liabilities does not seem to matter a lot. In addition, the non-core ratio correlates most tightly with cross-border liabilities vis-a-vis banks (89%), followed by liabilities vis-a-vis nonbank financials (80%) and all other sectors (17%). These correlations suggest that our non-core ratio captures to large extents banks’ interbank liabilities, an aspect that we will get back to in Section 5.2.
@@
Note. This figure plots the median non-core funding ratio across all of the banking systems in our sample, as well as separately for those located in countries with a fixed vs. floating exchange rate (evaluated with a one-quarter lag). The vertical dashed line marks the bankruptcy of Lehman Brothers in 2008:Q3. Sources: IFS, Ilzetzki et al. (2019).
@@ 2.2.1 Domestic Variables
Banking models—both those embedded in general equilibrium (Edwards and Vegh (1997), Catão and Rodriguez (2000), Céspedes et al. (2011)), as well as partial equilibrium ones (e.g., Bruno and Shin (2015))—show that increases in the (real) domestic interest rate and appreciations of the real exchange rate tend to favor foreign funding relative domestic-deposit funding, whereas liquidity and minimum capital requirements tend to depress lending and hence external borrowing by banks, all else constant. [^7] Considering a variety of regulatory instruments that include liquidity, loan-to-value and minimum capital rules, Alam et al. (2019) document sizeable effects on bank lending and private consumption growth in a cross-country panel dataset. To take into account the relevance of regulation, we control for (i) the country-level capital-to-asset ratio of banks and (ii) the tightness of macroprudential liquidity and capital regulation, based on the IMF’s integrated Macroprudential Policy (iMaPP) database, originally constructed by Alam et al. (2019). Specifically, the database reports for each country in the sample whether it tightens or loosens the degree of capital or liquidity regulation, respectively. As macroprudential regulation is likely to have a longer-lasting effect on bank balance sheets, we cumulate the two variables for each country.
@@
In order to study the extent to which the Mundellian trilemma—which postulates that it is impossible for a country to simultaneously have a stable exchange rate, full capital mobility, and an independent monetary policy (Mundell, 1963; Obstfeld et al., 2005)—affects the relation between the global financial cycle and bank balance sheets, we merge an exchange rate dummy equal to one for fixed exchange rate regimes (i.e., strictly fixed pegs), following the classification of Ilzetzki et al. (2019), to our dataset. Our sample is roughly balanced between countries with fixed and floating exchange rates, with 53% of the countries having a fixed exchange rate. As the average Chinn-Ito index (Chinn and Ito, 2006) in our sample stands at 0.93, our data set essentially consists of countries with open financial accounts only, and we hence do not explicitly control for this variable. [^8] Including this variable leaves our main results largely unchanged. This is chiefly an advantage of using a sample including only advanced economies, since, when examining the predictions of the Mundellian trilemma, one can focus on the role of the exchange rate regime. We winsorize all country-specific, independent variables at the 1% and 99% levels to limit the potential impact of outliers.
@@ 2.2.2 Global Variables
The other sub-set of explanatory variables comprises global factors. Those include fluctuations in risk aversion in global capital markets, world oil prices, the world non-bank share, weighted averages of inflation, GDP growth and macroprudential tightness of all of the economies in our sample, as well as the US shadow interest rate of Wu and Xia (2016), [^9] Relative to the US Federal Funds rate, this shadow rate has the advantage of not being bound at zero, so that the effects of QE can still be taken into account. and the dollar valuation relative to other currencies—the latter two reflecting the still hegemonic roles of the US financial system and of the US dollar in global finance and trade. Bruno and Shin (2015) as well as Coeurdacier et al. (2020), di Giovanni et al. (2021) and Obstfeld et al. (2019) underline the role of rising global risk aversion (as measured by the VIX) in dampening cross-border banking borrowing, hence potentially affecting the non-core to core ratio. In their survey of existing empirical work on the global financial cycle, Miranda-Agrippino and Rey (2022) emphasise that tighter US monetary policy tends to depress global asset prices and discourage global financial intermediation, at same time that they note that oil prices and other real global factors (proxied by global GDP and trade growth) can also play a role in international fluctuations in global bank credit (and hence on banks’ funding decisions). Other recent studies have emphasized the role of US dollar valuations in global capital flows and financial intermediation broadly defined (Avdjiev et al., 2019; Jiang et al., 2020), so it seems sensible to add a trade-weighted index of US dollar exchange rate as another potential global factor. We finally saturate the regressions with the worldwide average share of financial intermediation performed by nonbanks, assuming that a higher world non-bank share correlates with lower credit volumes and hence foreign borrowing of banks, as well as the GDP-weighted average of advanced economies’ cumulative macroprudential tightness, as we expect a tight link between global macroprudential tightness and banks’ funding choices. The latter variable is again based on the iMaPP Database. [^10] Breaking down this overall world macroprudential stringency index into the capital and liquidity sub-components leaves our results unchanged. Attendant results are available upon request. Table 1 provides the descriptive statistics for these global variables.
@@ 3 Empirical Approach
$y_{i,t}=\rho_{i}y_{i,t-1}+\boldsymbol{x}_{i,t}\boldsymbol{\beta}_{i}+\boldsymbol{g}_{t}\boldsymbol{\delta}_{i}+u_{i,t}$ (1)
$u_{i,t}=\boldsymbol{f}_{t}\boldsymbol{\gamma}_{i}+\epsilon_{i,t},$
where $y_{i,t}$ is the dependent variable and $y_{i,t-1}$ is its one-quarter lag. $\boldsymbol{x}_{i,t}$ is a $1\times K$ vector of strictly exogenous explanatory variables. $\boldsymbol{g}_{t}$ are the $M_{ocf}$ observed common factors, such as global oil prices, inflation and the VIX, and $\boldsymbol{f}_{t}$ are the $M_{ucf}$ unobserved common factors, modelled as a part of the overall unobserved multifactor error $u_{i,t}$ with the iid random noise term $\epsilon_{i,t}$. [^11] To avoid endogeneity of the right-hand side variables, we use one-quarter lags of all observable variables. Common factors affect all countries; however, due to the slope heterogeneity in the loadings $\boldsymbol{\delta}_{i}$ and $\boldsymbol{\gamma}_{i}$ the factors affect countries with a different strength. [^12] The additive fixed effects model is nested within the interactive fixed effects model if we assume the first element in $\boldsymbol{f}_{t}$ respectively the second element of $\boldsymbol{\gamma}_{i}$ to be one. The vector of unobserved coefficients $\boldsymbol{\beta}_{i}=[\beta_{i,1},...,\beta_{i,K}]^{\prime}$ is the parameter of interest and we allow it to vary across countries.
The advantage of the IFE model is that, in comparison to the widely used additive or two-way fixed effects model, it allows for the presence of unobserved common shocks, which are correlated with the explanatory variables. For example, assume that the country-specific vector of observed exogenous variables $\mathbf{x}_{i,t}$ is driven by the observed factors $\boldsymbol{G}_{t}$ and unobserved common factors $\boldsymbol{f}_{t}$:
$\mathbf{x}_{i,t}=\boldsymbol{g}_{t}\boldsymbol{\delta}_{i}^{x}+\boldsymbol{f}_{t}\boldsymbol{\gamma}_{i}^{x}+\xi_{i,t},$ (2)
then ignoring the observed and unobserved common factors will cause strong cross-sectional dependence, which can lead to biased and inconsistent estimates of $\beta_{i}$.
@@
If the common factors are correlated with the explanatory variables and omitted from the model, an omitted variable bias occurs. The most popular estimation methods in the literature to estimate models in the presence of strong cross-sectional dependence originating from unobserved common factors are the Common Correlated Effects (CCE) estimator (Pesaran, 2006) or Principal Components (PC) estimator (Bai, 2009, Moon and Weidner, 2015). The CCE estimator proved to be very flexible under different assumptions on the errors (Chudik et al., 2011), error structure (Kapetanios et al., 2011) or exogeneity (Chudik and Pesaran, 2015). Further, knowledge about the exact number of unobserved common factors is not required for the CCE estimator, which is an advantage over the PC estimator. The unobserved common factors are approximated by time-specific averages across all units, the so-called cross-sectional averages. The estimator is consistent if the number of unobserved common factors is smaller than the number of cross-section averages and if $N,T\rightarrow\infty$ with $T/N\rightarrow 0$. Denoting the dependent variable as the non-core to core ratio, the estimated equation using the CCE estimator becomes:
$y_{i,t}$ $=\rho_{i}y_{i,t-1}+\mathbf{x}_{i,t}\boldsymbol{\beta}_{i}+\boldsymbol{g}_{t}\boldsymbol{\delta}_{i}+\sum_{l=0}^{p}\bar{\boldsymbol{R}}_{t-l}\boldsymbol{\gamma}_{i,l}+e_{i,t}$ (3)
@@
where $\bar{\boldsymbol{R}}_{t}$ are the cross-section averages of the dependent and independent variables, excluding the observed common factors. Since Equation (3) includes lags of the dependent variable on the right-hand side, $p$ lags of the cross-section averages are added to approximate the unobserved common factors (Chudik and Pesaran, 2019). $e_{i,t}$ is the iid regression residual.
@@
Moving on to the specifics of our model, we distinguish the observed variables into country-specific variables and observed common factors:
@@
$\boldsymbol{x}_{i,t}$ $=\boldsymbol{\gamma}_{i}^{x}\boldsymbol{f}_{t}+\boldsymbol{\theta}_{i}^{x}\boldsymbol{g}_{t}+\boldsymbol{\chi}_{i,t}$ (7)
@@
The term $\boldsymbol{\chi}_{i,t}$ is neither driven by observed or unobserved common factors. $\boldsymbol{\omega}_{t}$ is the variation of the observed common factors that is independent of the unobserved common factors. Our main aim is the identification of those. However, only $\boldsymbol{x}_{i,t}$ and $\boldsymbol{g}_{t}$ are observed. Therefore ignoring the unobserved common factors will lead to an omitted variable bias.
@@
$y_{i,t}=$ $\rho_{i}y_{i,t-1}+\boldsymbol{\beta}_{i}\boldsymbol{\chi}_{i,t}+\left(\boldsymbol{\beta}_{i}\boldsymbol{\theta}_{i}^{x}+\boldsymbol{\theta}_{i}\right)\boldsymbol{\omega}_{t}$ (9)
$+\left[\boldsymbol{\beta}\left(\boldsymbol{\gamma}_{i}^{x}+\boldsymbol{\theta}_{i}^{x}\boldsymbol{\xi}\right)+\boldsymbol{\gamma}_{i}+\boldsymbol{\theta}_{i}\boldsymbol{\xi}+\epsilon_{i,t}\right].$
@@
The first two terms represent the effect of the country-specific variables and of the observed common factors on the dependent variable. The third term is the effect of the unobserved common factors. While the observed common factors are known and can be directly used, the unobserved common factors will be approximated by cross-section averages.
Our estimation approach consists of three steps. The first step is a CCE-MG estimation of (9) where the observed common factors are directly added and the cross-section averages span the remaining part of the factor space. An advantage of the CCE approach is that it incorporates the observed common factors in the same way as the observed country-specific variables. Therefore, as long as the number of cross-section averages exceeds the number of unobserved common factors, the coefficients $\boldsymbol{\beta}_{i}$ can be consistently estimated. The first step then approximates the unobserved common factors $\boldsymbol{f}_{t}$ with the cross-section averages, while the observed factors $\boldsymbol{g}_{t}$ are directly used. We then obtain an estimate of the composite term $u_{i,t}$ from Equation (5):

### v6

@@ 1 Introduction
Recent years have witnessed a heated debate in the macroeconomic and international finance literatures on the extent to which global financial conditions influence national ones and, within that, how far capital controls and flexible exchange rates can insulate national economies from global financial shocks. Following the seminal work of Rey (2015), as well as Borio (2014), Bruno and Shin (2013) and others, there has been widespread consensus that global financial conditions—epitomized by the concept of a global financial cycle—are a key determinant of domestic financial cycles and risk assessments by economic agents.
However, the jury is still out on the extent to which domestic policy instruments can insulate the domestic economy from the global financial cycle. Miranda-Agrippino and Rey (2020) and Passari and Rey (2015), using higher frequency data spanning a wider range of financial instruments over more than 30 years of data, have brought into question the effectiveness of domestic monetary policies and exchange rate flexibility as macroeconomic insulation devices, as advocated by Mundell (1963) in his famous depiction of the trilemma between exchange rate stability, monetary policy independence and free capital mobility. Yet, several papers, including Aizenman et al. (2008), Aizenman et al. (2016), Obstfeld et al. (2019), Obstfeld (2021) and Shambaugh (2004), find evidence that exchange rate flexibility can be effective in insulating domestic interest rates from the global financial cycle, at least partially. In addition, and over and above exchange rate flexibility, various recent studies find that financial account restrictions and some macro-prudential regulations can also be effective insulation devices (e.g., Akinci and Olmstead-Rumsey, 2018; Buch and Goldberg, 2018; Cerutti et al., 2018; Fendoğlu, 2017; Qureshi et al., 2011). These latter findings indicate that the Mundellian trillema remains, by and large, a suitable characterization of main trade-offs faced by domestic macroeconomic policy in a globalized world.
@@
As we discuss below, two important advantages of extracting principal components from residuals obtained by the CCE estimator are to allow (i) for the presence of unobserved or hard-to-measure global factors as well as (ii) for some cross-country heterogeneity in bank responses, which one might reasonably expect and which impart biases in the standard fixed effects estimator. We then establish how the choice of this approach influences our results relative to more standard panel fixed effects regressions in earlier studies. The end-goal is to establish the extent to which the estimated insulation role granted by exchange rate flexibility is sensitive to econometric estimation (or model) uncertainty.
@@
Our main findings are as follows. First, fluctuations in the ratio of non-core to core funding are highly persistent and overwhelmingly determined by global factors. Second, the combination of high cross-country dependence of non-core to core ratios and considerable heterogeneity in which global factors affect the banking system in each country, implies that the standard two-way fixed effects estimator (including country and time fixed effects) does not eliminate the cross-sectional dependence of the residuals, therefore producing likely inconsistent estimates of the response of bank ratios to the various country-specific and global disturbances. This indicates that the estimation methodology that we propose—combining the Pesaran (2006) CCE estimator to extract the relevant global factors (principal components) and then using the mean group estimator to tease out any differences between countries on a fixed vs. flexible exchange rate regime—should produce superior results. This estimation methodology then delivers our central result, namely that countries on a fixed exchange rate regime are more susceptible to have global factors affecting their bank funding ratios, or in other words exchange rate flexibility helps insulate the non-core to core ratio from such global factors. We also find that countries with tighter macroprudential policy stances are less affected by the global financial cycle, and this insulation is most effective for countries with a fixed exchange rate or in a monetary union. We finally take a closer look at the mechanisms through which the principal components affect banks’ non-core ratios. To this end, we first relate the principal components to the set of observed common factors, and find that most of the variation in banks’ non-core ratios can be explained by changes in the world macroprudential policy stance, the US short-term interest rate and financial crisis effects. In contrast, the VIX, oil prices, and the US real exchange rate generally have a smaller impact on banks’ non-core ratios. Then, using BIS Locational Banking Statistic data, we break down total foreign borrowing by banks into (i) domestic- vs. foreign-currency borrowing, as well as (ii) liabilities provided by foreign banks vs. foreign non-bank financials vs. other foreign sectors. In these specifications, we find that our results are driven mostly by banks’ interbank borrowing—a funding source that is known to be more flighty. The currency denomination does not seem to matter a lot.
@@ 2 Data
Our empirical analysis is based on a data set that combines two critical banking sector balance sheet variables from the International Financial Statistics at quarterly frequency over the period 2004:Q1-2022:Q1 with different domestic macroeconomic controls and global variables. Our sample includes the 31 advanced economies listed in Table A, [^5] In most specifications, we drop Iceland from our sample as it is an outlier for most variables. Its inclusion reduces the significance of our estimates, as we document in the robustness checks of Section 4.2. producing a balanced panel spanning more than 2,000 country-quarter observations. In this section, we give an overview about the data, with Appendix A providing all of the corresponding data details and sources.
@@ 2.1 Outcome Variables
Consistent with this evidence, we measure non-core funding ratios as the share of foreign liabilities over total deposits (IFS line 26c divided by IFS lines 24 and 25). To show how such a non-core ratio relates to other key balance sheet indicators highlighted in previous work, such as the ratio of bank loans-to-deposits (see Edwards and Vegh, 1997, Catão and Rodriguez, 2000, Obstfeld et al., 2005, 2019), we define loan-to-deposit ratio as IFS line 22d divided by IFS lines 24 and 25. As Table 2.1 shows, the average non-core ratio in our sample is equal to 78% and loan-to-deposit ratios have a mean of 122%. The variations are substantial with standard deviations of about 94% and 58%, respectively.
@@
Figure 1 shows that advanced economies’ non-core funding ratios are positively correlated ($\rho=0.92$) with swings in in banks’ loan-to-deposit ratios, revealing the important role of foreign borrowing in expanding lending beyond banks’ core funding sources (i.e., deposits). Figure 2 in turn shows that the non-core ratio is in general higher for countries on a fixed exchange rate. [^6] Notice that the floaters are in the mean on the left scale that is steadily two thirds lower than its span. Besides this level difference, it also shows conspicuous long-term trend shifts and considerable co-movement between the two country groups. In particular, non-core ratios increased significantly before the Global Financial Crisis, reaching a peak in 2008 where the global non-core ratio rose to 70%, to then decrease to a level of 30% in 2022. Only between 2015-20, there was some short-term divergence in the trends of fixers and floaters.
In the second part of the paper, we examine in more detail the mechanisms through which banks’ non-core funding ratios change. To this end, we employ data on banks’ cross-border liabilities from the BIS Locational Banking Statistics, which we also scale by total deposits. The data have poorer country coverage than our IFS data, but have the advantage of allowing for a breakdown into domestic- vs. foreign-currency liabilities, as well as into foreign liabilities provided by banks, nonbank financials, and all other sectors. The correlation between our benchmark non-core ratio and the overall BIS cross-border liability variable is 88%. Looking further into the different breakdowns of the BIS variable, we find that the correlation coefficient between the non-core ratio and the foreign-currency variable stands at 80%, whereas it is 82% for the domestic-currency variable, indicating that the currency denomination of foreign liabilities does not seem to matter a lot in our sample. In addition, the non-core ratio correlates most tightly with cross-border liabilities vis-a-vis banks (89%), followed by liabilities vis-a-vis nonbank financials (80%) and all other sectors (18%). These correlations suggest that our non-core ratio captures to large extents banks’ interbank liabilities, an aspect that we will get back to in Section 5.2.
@@
Note. This figure plots the median non-core funding ratio across all of the banking systems in our sample, as well as separately for those located in countries with a fixed vs. flexible exchange rate (evaluated with a one-quarter lag). The vertical dashed line marks the bankruptcy of Lehman Brothers in 2008:Q3. Sources: IFS, Ilzetzki et al. (2019).
@@ 2.2.1 Domestic Variables
Banking models—both those embedded in general equilibrium (Edwards and Vegh, 1997, Catão and Rodriguez, 2000, Céspedes et al., 2011), as well as partial equilibrium ones (e.g., Bruno and Shin, 2015)—show that increases in the (real) domestic interest rate and appreciations of the real exchange rate tend to favor foreign funding relative domestic-deposit funding, whereas liquidity and minimum capital requirements tend to depress lending and hence external borrowing by banks, all else constant. [^7] Considering a variety of regulatory instruments that include liquidity, loan-to-value and minimum capital rules, Alam et al. (2019) document sizeable effects on bank lending and private consumption growth in a cross-country panel dataset. To take into account the relevance of regulation, we control for (i) the country-level capital-to-asset ratio of banks and (ii) the tightness of macroprudential liquidity and capital regulation, based on the IMF’s integrated Macroprudential Policy (iMaPP) database, originally constructed by Alam et al. (2019). Specifically, the database reports for each country in the sample whether it tightens or loosens the degree of capital or liquidity regulation, respectively. As macroprudential regulation is likely to have a longer-lasting effect on bank balance sheets, we cumulate the two variables for each country.
@@
In order to study the extent to which the Mundellian trilemma—which postulates that it is impossible for a country to simultaneously have a stable exchange rate, full capital mobility, and an independent monetary policy (Mundell, 1963; Obstfeld et al., 2005)—affects the relation between the global financial cycle and bank balance sheets, we merge an exchange rate dummy equal to one for fixed exchange rate regimes (i.e., strictly fixed pegs), following the classification of Ilzetzki et al. (2019), to our dataset. Our sample is roughly balanced between countries with fixed and flexible exchange rates, with 53% of the countries having a fixed exchange rate, which includes membership of a monetary union. As the average Chinn-Ito index (Chinn and Ito, 2006) in our sample stands at 0.95, our data set essentially consists of countries with open financial accounts only, and we hence do not explicitly control for this variable. [^8] Including this variable leaves our main results largely unchanged. This is chiefly an advantage of using a sample including only advanced economies, since, when examining the predictions of the Mundellian trilemma, one can focus squarely on the role of the exchange rate regime. We winsorize all country-specific, independent variables at the 1% and 99% levels to limit the potential impact of outliers.
@@ 2.2.2 Global Variables
The other sub-set of explanatory variables comprises global factors. Those include fluctuations in risk aversion in global capital markets, world oil prices, the world non-bank share, weighted averages of inflation, GDP growth and macroprudential tightness of all of the economies in our sample, as well as the US shadow interest rate of Wu and Xia (2016), [^9] Relative to the US Federal Funds rate, this shadow rate has the advantage of not being bound at zero, so that the effects of QE can still be taken into account. and the dollar valuation relative to other currencies—the latter two reflecting the still hegemonic roles of the US financial system and of the US dollar in global finance and trade. Bruno and Shin (2015) as well as Coeurdacier et al. (2020), di Giovanni et al. (2021) and Obstfeld et al. (2019) underline the role of rising global risk aversion (as measured by the VIX) in dampening cross-border banking borrowing, hence potentially affecting the non-core to core ratio. In their survey of existing empirical work on the global financial cycle, Miranda-Agrippino and Rey (2022) emphasise that tighter US monetary policy tends to depress global asset prices and discourage global financial intermediation, with oil prices and other real global factors (proxied by global GDP and trade growth) also playing a role in international fluctuations in global bank credit (and hence banks’ funding decisions). Other recent studies have emphasized the role of US dollar valuations in global capital flows and financial intermediation broadly defined (Avdjiev et al., 2019; Jiang et al., 2020), so it seems sensible to add a trade-weighted index of the US dollar exchange rate as another potential global factor. We finally saturate the regressions with the worldwide average share of financial intermediation performed by nonbanks, assuming that a higher world non-bank share correlates with lower credit volumes and hence foreign borrowing of banks, as well as the GDP-weighted average of advanced economies’ cumulative macroprudential tightness, as we expect a tight link between global macroprudential tightness and banks’ funding choices. The latter variable is again based on the iMaPP Database. [^10] Breaking down this overall world macroprudential stringency index into the capital and liquidity sub-components leaves our results unchanged. Attendant results are available upon request. Table 2.1 provides the descriptive statistics for these global variables.
@@ 3 Empirical Approach
$y_{i,t}$ $=\rho_{i}y_{i,t-1}+\boldsymbol{x}_{i,t}\boldsymbol{\beta}_{i}+v_{i,t}$ (1)
$v_{i,t}$ $=\boldsymbol{f}_{t}\boldsymbol{\gamma}_{i}+\epsilon_{i,t},$ (2)
where $y_{i,t}$ is the dependent variable and $y_{i,t-1}$ is its one-quarter lag. $\boldsymbol{x}_{i,t}$ is a $1\times K$ vector of strictly exogenous explanatory variables. $\boldsymbol{f}_{t}$ are the $M_{ucf}$ unobserved common factors, modelled as a part of the overall unobserved multifactor error $v_{i,t}$ with the iid random noise term $\epsilon_{i,t}$. [^11] To avoid strong endogeneity of the right-hand side variables, we use one-quarter lags of all observable variables. Common factors affect all countries; however, due to the slope heterogeneity in the loadings $\boldsymbol{\gamma}_{i}$, the factors affect countries with a different strength. [^12] The additive fixed effects model is nested within the interactive fixed effects model if we assume the first element in $\boldsymbol{f}_{t}$ and for the two-way fixed effect model the second element of $\boldsymbol{\gamma}_{i}$ to be one. The vector of unobserved coefficients $\boldsymbol{\beta}_{i}=[\beta_{i,1},...,\beta_{i,K}]^{\prime}$ is allowed to vary across countries.
The advantage of the IFE model is that, in comparison to the widely used additive or two-way fixed effects model, it allows for the presence of unobserved common shocks, which can be correlated with the explanatory variables. For example, assume that the country-specific vector of observed exogenous variables $\mathbf{x}_{i,t}$ is driven by unobserved common factors $\boldsymbol{f}_{t}$:
$\mathbf{x}_{i,t}=\boldsymbol{f}_{t}\boldsymbol{\gamma}_{i}^{x}+\xi_{i,t},$ (3)
then ignoring the unobserved common factors will cause strong cross-sectional dependence, which can lead to biased and inconsistent estimates of $\mathbf{\beta}_{i}$.
@@
If the unobserved common factors are correlated with the explanatory variables and omitted from the model, an omitted variable bias occurs. The most popular estimation methods in the literature to estimate models in the presence of strong cross-sectional dependence originating from unobserved common factors are the Common Correlated Effects (CCE) estimator (Pesaran, 2006) or Principal Components (PC) estimator (Bai, 2009, Moon and Weidner, 2015). The CCE estimator proved to be very flexible under different assumptions on the errors (Chudik et al., 2011), error structure (Kapetanios et al., 2011) or exogeneity (Chudik and Pesaran, 2015). Further, knowledge about the exact number of unobserved common factors is not required for the CCE estimator, which is an advantage over the PC estimator. The unobserved common factors are approximated by time-specific averages across all units, the so-called cross-sectional averages. The estimator is consistent if the number of unobserved common factors is smaller than the number of cross-section averages and if $N,T\rightarrow\infty$ with $T/N\rightarrow 0$. The estimated equation using the CCE estimator becomes:
$y_{i,t}$ $=\rho_{i}y_{i,t-1}+\mathbf{x}_{i,t}\boldsymbol{\beta}_{i}+\sum_{l=0}^{p}\bar{\boldsymbol{R}}_{t-l}\boldsymbol{\gamma}_{i,l}+e_{i,t}$ (4)
@@
where $\bar{\boldsymbol{R}}_{t}$ are the cross-section averages of the dependent and independent variables. Since Equation (4) includes lags of the dependent variable on the right-hand side, $p$ lags of the cross-section averages are added to approximate the unobserved common factors (Chudik and Pesaran, 2019). $e_{i,t}$ is the iid regression residual.
@@
Moving on to the specifics of our model, denoting the noncore-to-core ratio as the dependent variable, we distinguish the observed variables into country-specific variables and $M_{ocf}$ observed common factors, $\boldsymbol{g}_{t}$, such as global oil prices, inflation and the VIX:
@@
$\boldsymbol{x}_{i,t}$ $=\boldsymbol{\gamma}_{i}^{x}\boldsymbol{f}_{t}+\boldsymbol{g}_{t}\boldsymbol{\theta}_{i}^{x}+\boldsymbol{\chi}_{i,t}$ (8)
@@
The term $\boldsymbol{\chi}_{i,t}$ is driven neither by observed nor unobserved common factors. $\boldsymbol{\omega}_{t}$ is the variation of the observed common factors that is independent of the unobserved common factors. Our main aim is the identification of $\boldsymbol{\chi}_{i,t}$ and $\boldsymbol{\omega}_{t}$. However, only $\boldsymbol{x}_{i,t}$ and $\boldsymbol{g}_{t}$ are observed.
@@
$y_{i,t}=$ $\rho_{i}y_{i,t-1}+\boldsymbol{\beta}_{i}\boldsymbol{\chi}_{i,t}+\boldsymbol{\omega}_{t}\left(\boldsymbol{\beta}_{i}\boldsymbol{\theta}_{i}^{x}+\boldsymbol{\theta}_{i}\right)$ (10)
$+\boldsymbol{f}_{t}\left[\boldsymbol{\beta}_{i}\left(\boldsymbol{\gamma}_{i}^{x}+\boldsymbol{\theta}_{i}^{x}\boldsymbol{\xi}\right)+\boldsymbol{\gamma}_{i}+\boldsymbol{\theta}_{i}\boldsymbol{\xi}\right]+\epsilon_{i,t}.$
@@
The first two terms represent the effect of the country-specific variables and of the observed common factors on the dependent variable. The third term is the effect of the unobserved common factors. Although the observed common factors are known and can be directly used, the unobserved common factors will be approximated by cross-section averages. We further define a composite, $u_{i,t}$, which contains the observed and unobserved common factors: